OpenStudy (rebecca1588):

Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates: 2 4 6 8 Cost: 25 45 65 85 A. y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) D. y – 85 = 10(x – 8)

OpenStudy (rebecca1588):

I think its D or B, I could use a second opinion

OpenStudy (rebecca1588):

@superdavesuper I think its D or B

OpenStudy (anonymous):

If we try and write it like an equation, we can call 'crates' our x value and 'cost' our y value. Take any two points from your data set e.g. (2, 25) and (4,45) Now slope (m) = change in y / change in x m = (45-25)/(4-2) m = 20/2 m = 10 We can write equations as (y-b) = m(x-a) where m is the slope and (a,b) is any point (y-85) = 10(x-8) D

OpenStudy (rebecca1588):

Thanks so much <3 :)

OpenStudy (anonymous):

You're welcome :)